OpenStudy (algtrigcalc):

Can someone show me how to find the probability algebraically? Two dice are rolled. Find the probability of each event. a. the sum of the number is 10. b. the sum is at least 10. c. at least one die shows a 4. d. exactly one die shows a 4.

jimthompson5910 (jim_thompson5910):

Use a dice chart like this

jimthompson5910 (jim_thompson5910):

focus on the numbers in black how many numbers are 10?

OpenStudy (algtrigcalc):

3 numbers

jimthompson5910 (jim_thompson5910):

out of how many black numbers total?

OpenStudy (algtrigcalc):

3/36 = 1/12

jimthompson5910 (jim_thompson5910):

very good, that's the final answer to part A

OpenStudy (algtrigcalc):

Is there a way to do it algebraically? In case my teacher gives me larger numbers?

jimthompson5910 (jim_thompson5910):

well there are 6 ways to roll 1 die, so there are 6*6 = 36 ways to roll both dice

jimthompson5910 (jim_thompson5910):

there are 3 ways to add to 10 4+6 = 10 5+5 = 10 6+4 = 10

jimthompson5910 (jim_thompson5910):

There isn't an algebraic way in terms of having a formula or solving for x or anything like that.

OpenStudy (algtrigcalc):

Oh ok

OpenStudy (algtrigcalc):

I understand the rest. Thank you!

jimthompson5910 (jim_thompson5910):

I'm glad I could help